### v1

# Distributed Split Computing Using Diffusive Metrics for UAV Swarms

Abstract
In large-scale UAV swarms, dynamically executing machine learning tasks can pose significant challenges due to network volatility and the heterogeneous resource constraints of each UAV. Traditional approaches often rely on centralized orchestration to partition tasks among nodes. However, these methods struggle with communication bottlenecks, latency, and reliability when the swarm grows or the topology shifts rapidly. To overcome these limitations, we propose a fully distributed, diffusive metric-based approach for split computing in UAV swarms. Our solution introduces a new iterative measure, termed the aggregated gigaflops, capturing each node’s own computing capacity along with that of its neighbors without requiring global network knowledge. By forwarding partial inferences intelligently to underutilized nodes, we achieve improved task throughput, lower latency, and enhanced energy efficiency. Further, to handle sudden workload surges and rapidly changing node conditions, we incorporate an early-exit mechanism that can adapt the inference pathway on-the-fly. Extensive simulations demonstrate that our approach significantly outperforms baseline strategies across multiple performance indices, including latency, fairness, and energy consumption. These results highlight the feasibility of large-scale distributed intelligence in UAV swarms and provide a blueprint for deploying robust, scalable ML services in diverse aerial networks.
Index Terms:
Split Computing, UAV Swarms, Distributed Optimization, Diffusive Metrics
I Introduction
The escalating prominence of Unmanned Aerial Vehicle (UAV) swarms has opened up new horizons in applications spanning surveillance, disaster response, smart agriculture, and beyond. By coordinating large numbers of autonomous aerial units, system-level benefits such as enhanced coverage, redundancy, and adaptability can be realized. However, these advantages come hand-in-hand with steep technical challenges, particularly when it comes to designing and deploying machine learning (ML) services in vast swarms. Traditional methods that rely on a single, centralized server or controller become increasingly untenable as the number of nodes and geographic scale increase[1]. A centralized ground station, faces prohibitive communication overheads, latency bottlenecks, and reliability issues. Moreover, real-world UAV swarms are dynamic, meaning that nodes may join, leave, or rearrange themselves in the network. Such dynamic conditions only make the limitations of traditional centralized paradigms worse[2].
Recent efforts to alleviate centralized burdens typically advocate a form of distributed operation by distributing inference or training tasks across multiple nodes[3]. However, they still often rely on global awareness or explicit synchronization. At smaller scales or in constrained network topologies, these partial-distribution schemes might suffice. Yet, as UAV swarms expand to hundreds or even thousands of nodes, it becomes infeasible to gather and coordinate global network information[4]. Additionally, heterogeneous node capabilities such as differences in on-board processing, sensing, and power resources demand an approach that can dynamically adjust to varying node states in near real-time. A node endowed with high computation power a time may deplete its battery and become less capable a moment later, or an agile UAV that was ideally placed to handle tasks at one moment could relocate out of range in the next. These variations highlight the necessity for a self-organizing approach that pushes intelligence and decision-making to the edges.
Split learning and computing is particularly well-suited to address these challenges in large-scale UAV swarms[3]. Rather than transmitting raw data or entire models to a single aggregator, split learning and computing partitions the model into layers distributed among various nodes, thereby enabling partial inference at local levels. By transferring only the necessary, intermediate representations, nodes can reduce communication overhead while maintaining data privacy. However, effectively implementing split computing in UAV networks requires strategic decision-making about when and where to forward intermediate layer outputs[1, 3]. This is where centralized schemes fail to perform well for large networks. They attempt to collect global knowledge of the network’s topology, node resource levels, and data distributions which is impractical in large, dynamic swarms. As a result, model splits and partial inferences often end up bottlenecked by nodes that unknowingly become overloaded or unreachable.
To overcome these limitations, we propose a fully distributed, diffusive metric-based approach in which each node autonomously assesses its own capabilities as well as those of its neighbors without relying on global information. The core idea is to estimate the “aggregated gigaflops” of a node in a manner analogous to graph centrality metrics. In particular, our approach treats each node’s local computational power as well as its neighbors’ powers in an iterative fashion: if a node’s neighbors can handle substantial partial inferences, or if they hold a topologically advantageous position to further distribute the load, the node is more inclined to offload its tasks there. By repeating local updates, the entire network reaches a stable state of partial inference routing in which tasks flow toward underutilized, resource-rich parts of the swarm. This distributed decision-making ultimately fosters better throughput, lower latency, and more balanced energy consumption. Our method thereby aligns with the inherent nature of UAV swarms: distributed, dynamic, and operating over sprawling distances that defeat naive centralized solutions.
To further enhance the system’s adaptability under fluctuating workloads, we incorporate an early-exit mechanism that dynamically reduces the depth of inference based on real-time load conditions. This strategy allows nodes to truncate the model execution at predefined intermediate layers, producing acceptable outputs while avoiding the full computational cost of processing the entire model. By adjusting the exit point according to the current resource availability and task queue length, the system minimizes latency and prevents resource saturation. This approach aligns with the principles of early exiting [5] where intermediate outputs offer a controllable trade-off between computational effort and inference accuracy, making it particularly suitable for dynamic and resource-constrained UAV environments.
By merging the advantages of split computing with a diffusive, graph-oriented update rule for computational metrics, our approach maintains robust inference quality across a large and rapidly evolving topology. It enables UAV swarms to shift inferences efficiently among nodes, ensuring that tasks do not pile up in isolated regions while idle capacity elsewhere remains unused. In contrast with conventional centralized strategies, which suffer from communication bottlenecks and single points of failure, our framework supports agility and scalability by design.
In this paper, our key contributions are as follows:
  • We identify the fundamental scalability challenges in deploying split computing across large-scale, dynamic UAV swarms, highlighting why centralized solutions are prone to failure in real-world scenarios.
  • We introduce a novel metric, the aggregated gigaflops, to capture how a node’s computational load should be shared in conjunction with its neighbors’ resources. This metric is designed to be iteratively updated at each node, enabling a fully distributed and lightweight approach.
  • We propose a new decision-making framework for routing partial inferences (either intermediate feature maps or raw input data) based on the aggregated gigaflops metric. This framework avoids reliance on global synchronization, making it highly scalable and robust to node mobility or failure.
  • We design an early-exit mechanism that dynamically truncates model execution based on local load conditions, reducing latency and preserving system stability under high workload scenarios.
  • We conduct extensive simulations to verify that our distributed solution outperforms existing benchmark strategies across several metrics including inference performance, energy efficiency, fairness, and overall throughput, particularly in large UAV swarms.
The remainder of this paper is organized as follows. Section II reviews related work on distributed inference and task offloading in UAV networks. Section III formalizes the system model and problem definition, introducing the proposed diffusive metric and task allocation strategy. Section V presents the simulation setup and provides a detailed performance evaluation under various operating conditions. Finally, Section VI concludes the paper and outlines directions for future work.
Figure 1: Proposed distributed Split-Learning that uses diffusive aggregated Gflops.
II Related Work
Recent research has explored distributed DNN inference and task offloading to address the challenges posed by limited resources and dynamic network conditions. In [6], MoDNN partitions DNN computations among mobile devices to balance workload and reduce communication overhead; its model partitioning strategy minimizes transmission delays while exploiting parallel processing, making it highly applicable to resource-constrained environments. Similarly, DISNET [7] introduces a micro-split framework that partitions deep neural network models both vertically and horizontally, enabling cooperative inference among heterogeneous IoT devices while reducing latency and energy consumption through adaptive workload allocation.
On the split computing front, [8] propose a multi tier framework that employs deep reinforcement learning to optimize model splitting and resource allocation in UE–Edge–Cloud systems, decomposing complex optimization problems into sequential subproblems to balance computational and communication loads. In a similar vein, CoEdge [9] demonstrates that dynamic workload partitioning—formulated as a constrained optimization problem—can yield substantial latency speedups, with a fast approximation algorithm that adapts in real time to varying device capabilities and network conditions. Similarly for Split Learning, [10] develop an efficient parallel split learning (EPSL) framework for resource-constrained wireless edge networks. Their approach jointly optimizes subchannel allocation, transmit power, and model split point selection to minimize latency and communication overhead. [11] propose Edge-MSL, a split learning framework for mobile edge networks that leverages a contextual multi-armed bandit formulation to dynamically select the optimal split point and assign edge servers. By adapting the split location and offloading decisions based on real-time network conditions, their approach significantly reduces training latency and enhances convergence. On the other hand, [12] combines split learning with federated learning in UAV networks to address the trade-off between communication overhead and computational constraints, resulting in an architecture that achieves higher inference accuracy than standalone federated learning while reducing transmission costs compared to conventional split learning. Moreover, [13] propose a parallel split learning framework over wireless networks that integrates dynamic resource management with parallel task offloading. Their method exploits concurrent processing across heterogeneous edge devices to substantially reduce training latency while maintaining balanced workload distribution. [14] investigate a UAV-assisted split computing system where a UAV acts as a mobile edge server to support collaborative inference among IoT devices. By partitioning a DNN between on-device processing and UAV-based computation, their design reduces transmission delays and energy consumption.
Task offloading approaches can also be useful for split learning approaches. For instance, [15] develops a UAV-based environmental monitoring system that employs on-board processing to minimize transmission delays and maintain low data freshness latency, thereby enhancing overall system capacity in remote areas. [16] proposes a cloud–MEC collaborative offloading scheme that integrates differentiated offloading decisions with service orchestration via an ODaS mechanism, significantly reducing both delay and energy consumption by balancing computational and communication loads. [17] addresses the challenge of offloading dependent tasks modeled as directed acyclic graphs using a branch soft actor-critic algorithm, which optimizes offloading, computation, and cooperation decisions under dynamic conditions to reduce latency and improve energy efficiency. More recently, [18] presents a multi-agent collaborative framework that jointly optimizes UAV trajectory planning and DAG task offloading via deep reinforcement learning, with their TD3-TT algorithm demonstrating further latency reductions and enhanced real-time performance. The FOCUS framework [19] leverages fog computing in UAV networks by jointly optimizing routing and computation. Its central SDN-based architecture and analytical approach achieve significant latency improvements by effectively integrating on-board processing with data forwarding. Lastly, [20] investigate task offloading in a UAV-assisted MEC network by leveraging a federated learning framework. In their work, a UAV equipped with MEC capabilities serves as a mobile server to assist resource-constrained UE in offloading computation-intensive tasks.
Distinct from these approaches, our work targets a fully distributed split computing framework with early-exit mechanism, eliminating the need for centralized coordination in large-scale UAV swarms.
III Problem Formulation and System Model
Figure 1 provides an overview of our proposed distributed split computing framework within a large and dynamic UAV network. The core challenge lies in determining how and where to execute computation tasks to achieve low latency, balanced resource utilization, and robust performance, especially when the network topology changes frequently and UAVs have heterogeneous capabilities. We assume that the full neural network model, along with its corresponding early-exit parts, is pre-trained and deployed on the UAVs prior to operation. This ensures that each node can execute both complete and truncated inferences locally, according to its current workload and resource availability. To address these challenges, our formulation focuses on four interlinked aspects: (i) modeling communications among UAVs, (ii) defining a distributed metric for aggregated computation capability, (iii) making task-transfer decisions based on local utilization, and (iv) applying an early-exit mechanism to adapt to real-time workload fluctuations. The following subsections describe each component in detail.
III-A Task Computation Model
We consider machine learning tasks that are partitioned into L layers, represented as nodes in a DAG for the Machine learning task. The execution order of this task $T$ is decided by this DAG. Each layer $l\in\{1,2,\ldots,L\}$ requires $G_{l}$ gigaflops to complete. Each UAV $v_{i}$ provides a computational capacity of $F_{i}$ gigaflops per second. Thus, the delay to compute layer $l$ at UAV $v_{i}$ is given by
$d^{\mathrm{comp}}_{l,i}=\frac{G_{l}}{F_{i}}$ (1)
When processed entirely on a single node $v_{i}$, the overall computation delay is determined by the sum of delays along the longest dependency chain in the DAG. For tasks with purely sequential dependencies, this simplifies to
$d^{\mathrm{comp}}_{T,i}=\sum_{l=1}^{L}d^{\mathrm{comp}}_{l,i}$ (2)
In scenarios where intermediate layer outputs are offloaded during processing, if the offloading decision occurs before the completion of a computation, any partially computed work is discarded.
III-B Communications
We consider a UAV network in which nodes are distributed across a region and follow circular trajectories. Each UAV communicates using a two-ray ground-reflection model, allowing us to calculate the channel capacity between nodes $v_{i}$ and $v_{j}$ via Shannon’s formula:
$C_{ij}=B\log_{2}(1+SNR_{ij})$ (3)
The term $SNR_{ij}$ is signal to noise ratio of the received transmission from node $v_{i}$ to node $v_{j}$ and B is the bandwidth of the channel in Hertz. Calculation of $SNR_{ij}$, needs pathloss to be known. For this, we calculate pathloss using two ray ground propagation model, where receiver is assumed to receive only one other copy of the original transmission reflected from the ground. Then, signal to noise ratio for transmission from node $v_{i}$ to node $v_{j}$, $SNR_{ij}$ can be calculated as:
$SNR_{ij}=P_{i}+L(v_{i},v_{j})-N_{0}$ (4)
Here, $P_{i}$ is the transmit power of the node $v_{i}$ and, $N_{0}$ is the mean of the thermal noise. Lastly, $L(v_{i},v_{j})$ is the pathloss function that calculates the pathloss between nodes $v_{i}$ and $v_{j}$.
III-C Energy Consumption Model
For energy consumption, we model that every gigaflops computed consumes $J$ joules. Therefore, the energy required to compute unit $l$ is
$E_{l}=G_{l}\times J$ (5)
and the total energy consumption for a task processed entirely on one node is
$E_{T}=\sum_{l=1}^{L}G_{l}\times J$ (6)
Additionally, when offloading a computational unit from one UAV to another, the transmission energy consumption is given by:
$E^{tx}_{l}=\frac{S_{l}}{C_{ij}}\times P_{i}$ (7)
where $S_{l}$ denotes the data size of output of layer $l$, $C_{ij}$ is the channel capacity between nodes $v_{i}$ and $v_{j}$, and $P_{i}$ is the transmission power of node $v_{i}$. This expression represents the product of the transmission time $\left(\frac{S_{l}}{C_{ij}}\right)$ and the power used during transmission.
IV Distributed Task Allocation and Early-Exit Mechanism
We adopt a distributed approach for UAVs because the network spans a large geographic area with dynamic UAVs. Centralizing the network would lead to significant communication overhead. Our decision-making focuses on offloading the current machine learning computation to the least utilized node. Utilization is determined by the aggregated computing capability of nodes, considering the resources of neighboring nodes.
Nodes with powerful neighbors or central positions are more advantageous for task allocation. To identify such nodes, we aim to utilize graph centrality measures. However, these measures typically require a global view of the network, which conflicts with our distributed approach. Metrics like eigenvector centrality, which consider only a node’s neighbors, provide a feasible alternative. Eigenvector centrality for a node is calculated as:
$x_{v_{i}}={\frac{1}{\lambda}}\sum_{v_{j}\in M(v_{i})}x_{v_{j}}$ (8)
Here $M(v_{i})$ is the set of neighbors of $v_{i}$ and $\lambda$ is the largest eigenvalue of the network graph’s adjacency matrix. Eigenvector centrality is iteratively computed using power iteration: initial centralities are guessed and updated until convergence. Normalization is required at each step to prevent divergence, though this typically requires a global view. This iterative approach could work for distributed networks if divergence can be controlled.
To address these challenges, we propose using aggregated computation capability instead of eigenvector centrality. This metric measures how quickly a node can complete a job when it distributes the load among itself and its neighbors.
Figure 2: Proposed early-exit mechanism
We iteratively compute aggregated computation capability. The idea is if a load given to a node, if it were to distribute this load to its neighbors evenly, when this load would be completed. The aggregated computation capability $\phi_{i,t+1}$ of node $v_{i}$ at time $t+1$ is:
$\frac{1}{\phi_{i,t+1}}=\frac{1}{|M_{i}(t)|+1}\big{(}\frac{1}{F_{i}}+\max_{k\in M_{i}(t)}\{d^{tx}_{i,k}+\frac{1}{{\phi_{k,t}}}\}\big{)}$ (9)
where $M_{i}(t)$ is the neighbor set of node $i$, $F_{i}$ is computation capability of node $i$, $d^{tx}_{i,k}$ is the transmission delay between nodes $i$ and $k$, and $\phi_{i,t}$ is the aggregated computation capability of node $i$ and time $t$. The aggregated computation capability assumes that the node divides the work evenly ($|M(i)|+1$), computes its part, then it broadcasts the load to its neighbors. Since the neighbors can work concurrently, the task will be completed lastly by the slowest node with aggregated computation capability. Since this update requires only the neighbors of the nodes, it dynamically adapts the network conditions without need for global view.
IV-A Task Transfer Decision
Our framework requires each UAV to make autonomous decisions about whether to process a task locally or transfer it to a neighboring node. These decisions are based on comparative utilization levels between nodes, mediated by a threshold parameter $\gamma$ that controls the sensitivity of task offloading.
The utilization of node $i$ at time $t$ is calculated as the ratio of its current computational load to its aggregated computation capability:
$U_{i}(t)=\frac{T_{i,t}}{\phi_{i,t}}$ (10)
where $T_{i,t}$ represents the total computational load in the task queue of node $i$ at time $t$ measured in gigaflops, and $\phi_{i,t}$ is the node’s aggregated computation capability.
When deciding whether to transfer a task, node $i$ first identifies the neighboring node with minimum utilization:
$k^{*}=\underset{k\in M_{i}(t)}{\arg\min}\ U_{k}(t)$ (11)
The task transfer decision is then governed by a piecewise function using the threshold parameter $\gamma$:
$\delta_{i\rightarrow k^{*}}(t)=\begin{cases}\text{Transfer to node }k^{*},&\text{if }U_{i}(t)-U_{k^{*}}(t)>\gamma\\ \text{Process locally},&\text{otherwise}\end{cases}$ (12)
The parameter $\gamma$ serves as a stability threshold that prevents excessive task transfers when utilization differences are minimal. A higher value of $\gamma$ makes the system more conservative, requiring a larger utilization gap to justify offloading, which reduces communication overhead but may lead to less balanced workloads. Conversely, a lower $\gamma$ makes the system more sensitive to utilization differences, promoting load balancing at the potential cost of increased communication overhead.
This mechanism ensures that tasks are only transferred when the utilization difference exceeds a meaningful threshold, preventing oscillatory behavior where tasks might be repeatedly transferred between nodes with similar utilization levels.
IV-B Early-Exit Mechanism
To maintain consistent performance under varying workloads, nodes dynamically adjust the depth of the model inference based on real-time buffer conditions. This mechanism integrates an early-exit strategy that reduces computational load and latency by truncating the network execution at predefined layers, thereby producing intermediate outputs with acceptable accuracy levels. Figure 2 illustrates the proposed early-exit mechanism. Each node monitors the rate of change in its task queue by computing the time-normalized derivative of the remaining task GFLOPs:
$\Delta T_{i,t}=\frac{T_{i,t}-T_{i,t-1}}{\Delta t},$ (13)
where $\Delta T_{i,t}$ is the first difference of the remaining task GFLOPs and $\Delta t$ is the interval between allocation steps. To account for abrupt fluctuations (e.g., due to new task creation or completion), we smooth this difference:
$D_{i,t}=D_{i,t}+\alpha(\Delta T_{i,t}-D_{i,t}),$ (14)
where $D_{i,t}$ is the smoothed derivative and $\alpha$ controls the smoothing factor. A high $D_{i,t}$ value indicates sustained task accumulation, signaling that node $v_{i}$ is unable to process its workload efficiently.
To prevent overload and reduce inference latency, each node adjusts its inference depth based on the smoothed derivative $D_{i,t}$ of its task queue. Let $\tau_{\mathrm{med}}$ and $\tau_{\mathrm{high}}$ denote the medium and high congestion thresholds, respectively. The early-exit decision is characterized by the exit label $\xi_{i,t}$, defined as:
$\xi_{i,t}=\begin{cases}L_{\text{full}},&\text{if }D_{i,t}\leq\tau_{\mathrm{med}},\\ L_{1},&\text{if }\tau_{\mathrm{med}}<D_{i,t}\leq\tau_{\mathrm{high}},\\ L_{2},&\text{if }D_{i,t}>\tau_{\mathrm{high}},\end{cases}$ (15)
where $L_{\text{full}}$ indicates processing the full network, $L_{1}$ denotes truncation layer when there is medium congestion, and $L_{2}$ is for the high congestion. In both early-exit cases ($L_{1}$ and $L_{2}$), an additional three layers are processed after the exit point to finalize the output.
This adaptive mechanism ensures that nodes under high load degrade computation gracefully, trading off accuracy to maintain responsiveness and system stability. It is important to note that the neural network architecture, the selected split points, and the number of layers used in this work serve as a specific example corresponding to the evaluation presented in Section V. In practice, the neural network designer must analyze the target model to identify appropriate split points and determine the number of additional layers to execute after each truncation. These decisions depend on the network’s structure and the desired balance between accuracy and computational efficiency under dynamic conditions.
V Performance Evaluation
Our simulation environment is designed to emulate the dynamic conditions of a large-scale UAV swarm executing split computing tasks. In our environment, UAVs are deployed over a 20×20 km area and follow circular trajectories with a movement radius of 1000 m, traveling at speeds of up to 75 m/s. Task arrivals are modeled as a Markov process with an average inter-arrival time of 60 ms, which accurately captures the random and bursty nature of workload generation. In addition, the simulator captures detailed communication and computation aspects—including realistic channel characteristics, transmission delays, and the heterogeneous computational capabilities of each UAV ensuring that both the physical movement and the fluctuating workload distribution among nodes are properly considered. Table I shows the rest of the simulation parameters.
Within this dynamic simulation framework, we evaluate five distinct task offloading strategies. In the first scenario, tasks are offloaded in a purely random manner to a neighboring UAV, providing a basic baseline for performance comparison. The second strategy follows a similar random selection process, but enforces an acyclic constraint to prevent redundant offloading cycles. The third approach adopts a greedy policy, where each UAV offloads its task to the least busy neighbor based on instantaneous load measurements, thus aiming to minimize processing delays. The fourth strategy, referred to as LocalOnly, eliminates offloading entirely so that tasks are processed solely on the originating UAV, which highlights the trade-off between local processing overhead and the benefits of collaborative offloading. Finally, our proposed Distributed method leverages a diffusive metric—in particular, the aggregated gigaflops measure—to dynamically route tasks based on both the local UAV’s capacity and the capabilities of its neighbors. This comprehensive set of cases allows us to rigorously assess how different offloading mechanisms perform under the challenging conditions of a mobile, Markov-driven UAV swarm.
In our experiments, we consider the neural network model described in Section III pre-trained offline and deployed across the UAV swarm. The full model achieves an accuracy of 0.95, while the early-exit configurations at the 30th and 15th layers provide accuracies of 0.9 and 0.6, respectively. These values quantify the trade-off between computational savings and inference quality when applying early exits under high-load conditions. Lastly, the threshold $\gamma$ for distributed allocation is set to 0.02 based on our empirical data gathered from the empirical simulations. We selected this value to effectively balance the trade-off between frequent task transfers and achieving equitable workload distribution.
TABLE I: Simulation Parameters
Parameter Value
Number of Workers 10-50
Area Bounds 20x20 km
Node Placement Granularity 15
Movement Radius 1000 m
Movement Speed 75 m/s
Worker Capability $(\mu,\sigma)$ N(400,100) GFlops
Energy per GFlop 0.02 J / GFlop
Task Generation Period 60 ms
Exit Points $(L_{\text{full}},L_{1},L_{2})$ [15, 30, 60]
Exit Thresholds $(\tau_{\mathrm{med}},\tau_{\mathrm{high}})$ [1.5, 2.5]
Exit Accuracy Levels [0.6, 0.9, 0.95]
Transmission Power 30 dBm
Noise Power -85 dBm
Minimum SNR 3 dB
Bandwidth 10 MHz
Number of Runs 50
Maximum Simulation Time 100 s
Random Task Allocation Probability 0.2
Random Acyclic Allocation Probability 0.1
Distributed Allocation $\gamma$ 0.02
Decision Period 200 ms
Greedy Allocation Probability 0.05
V-A Basic Simulations
The initial set of simulations examines a baseline scenario that does not employ early-exit mechanism. Figure 3(a) illustrates the average task latency as a function of the number of workers, demonstrating that latency decreases as more workers are available. In Figure 3(b), the average remaining GFLOPs in UAV buffers are depicted, providing insight into workload accumulation and balancing among nodes. Figure 3(c) presents the average transfer times, where the greedy approach minimizes transfers initially but experiences an increase as load builds, while the distributed method transfers only when necessary. Jain’s fairness index is reported in Figure 3(d) by normalizing the processed GFLOPs with respect to each UAV’s capability, thereby ensuring that more capable nodes receive a proportionally higher workload. Figure 3(e) shows the energy consumption per task, where the local-only method benefits from avoiding transmission overhead.
Lastly, we come up with a figure of merit (FOM) that combines previously measured metrics to gauge the overall performance of the methods in (16). The FOM is measured in Figure 3(f) and illustrates that our proposed distributed method consistently achieves the highest performance across all worker counts, although the local-only approach remains competitive at lower worker densities due to a reduced number of generated tasks.
$\text{FOM}=\frac{\text{TPS}\times\text{ACC}}{\text{AE}\times\text{AL}}$ (16)
where TPS is task completed per second, ACC is average accuracy of completed tasks, AE is average energy usage, and AL is average latency per task.
Overall, the results presented in Figures 3 collectively demonstrate the effectiveness of the proposed distributed method in balancing key performance trade-offs across varying swarm sizes. Specifically, the distributed approach achieves consistently lower latency, reduced workload accumulation, and minimal transfer overhead compared to alternative strategies. Moreover, it maintains a high degree of fairness in task distribution and favorable energy consumption, despite the added complexity of inter-node coordination. These advantages become more pronounced as the number of workers increases, highlighting the scalability of the proposed solution. The superiority of the distributed method is further confirmed by the FOM index, which integrates all core performance metrics into a single indicator and consistently favors the proposed scheme across the entire range of worker densities. These findings underscore the capability of our system to efficiently manage computational loads in dynamic UAV environments without centralized control, while ensuring both high resource utilization and robustness against congestion.
Figure 3: Simulation results as a function of the number of workers. (a) shows average latency, (b) depicts remaining GFLOPs, (c) illustrates average transfer time, (d) presents Jain’s fairness index, (e) reports energy per task, and (f) shows the overall figure of merit.
V-B Effect of Input Rate
Next, we measure the effect of input rate when there is fixed number of workers. We vary mean task arrival time from 60ms to 100ms for 30 workers. Figure 4(a) has the average task latency results. Our proposed method outperforms others for high task input rates (lower input periods). However as task arrival time increases, the gap between LocalOnly and out method diminishes. Because, at this point most of the edge nodes are idle and are not overloaded. Thus, overhead of transmitting tasks actually hurt the performance as in Random methods. As a result, our method can not find opportunities to collaborate which makes it perform similar to LocalOnly. Moreover, in figure 4(b) we take a look into remaining task GFLOPs in edge node buffers. The behavior is similar to latency results. One thing to notice here is that lower latency does not result in lower remaining task GFLOPs. Because a task can be half processed, have less remaining GFLOPs and still completed later. The difference of LocalOnly and Random methods illustrate this clearly. Lastly, we measure the FOM for all methods in figure 4(c). When the network is stressed, the LocalOnly method can not keep up with Greedy and our Distributed method. One interesting result here is that, when the network is relaxed the LocalOnly is the overall best method. This makes sense because when there are not many tasks, the overhead of collaboration does provide significant benefits.
Figure 4: Performance metrics under varying task arrival rates for 30 workers. Panel (a) shows average latency, (b) displays remaining GFLOPs, and (c) presents the overall figure of merit.
V-C Effect of Network Distribution
The performance of collaborative task offloading is further influenced by the spatial distribution of UAVs, as captured by variations in the mission area size. Figure 5(a) shows that as the mission area increases, the opportunities for inter-node collaboration decline, resulting in increased overall latency. This effect is particularly pronounced in larger areas where the network connectivity weakens, causing the system to behave more like a collection of isolated local processors rather than an integrated distributed system. The impact on workload balancing is further illustrated in Figure 5(b), where larger mission areas correlate with a higher average of remaining GFLOPs in certain nodes, indicating that the collaborative mechanism is less effective in redistributing the load uniformly. Consequently, the overall figure of merit (Figure 5(c)) trends downward as the mission area expands, underscoring the sensitivity of the distributed strategy to network density and connectivity. These observations suggest that for large-scale deployments, additional strategies to enhance connectivity or adaptive routing protocols may be necessary to sustain optimal performance.
Figure 5: Performance metrics under varying mission area sizes. Panel (a) shows average latency, (b) depicts remaining GFLOPs, and (c) presents the overall figure of merit.
V-D Early-Exit Mechanism
In the final set of experiments, we evaluate the efficacy of our early-exit mechanism, which adapts the network configuration in real-time based on load and buffer conditions executing an early-exit mechanism. This mechanism is designed to preemptively mitigate bottlenecks by allowing overloaded nodes to reduce their local computation and offload tasks to less burdened neighbors. The benefits of this approach are clearly evident across multiple metrics. For instance, Figure 6(a) demonstrates that early-exit helps maintain high inference accuracy by preventing nodes from becoming overwhelmed. Similarly, the average latency, as shown in Figure 6(b), is significantly reduced when early-exit is enabled. The reduction in latency can be attributed to the more efficient distribution of computational tasks, which minimizes the time that tasks spend waiting in overloaded buffers.
Figure 6(c) provides insights into the distribution of remaining GFLOPs across the network, highlighting that early-exit results in a more even distribution of workload. This balance is further corroborated by the fairness index shown in Figure 6(d), which indicates that the distributed processing is more equitably spread among the UAVs. Energy efficiency, an important consideration in UAV operations, is also enhanced by early-exit mechanism, as evidenced by the lower average energy per task depicted in Figure 6(e). Ultimately, the overall figure of merit (Figure 6(f)) encapsulates these improvements, demonstrating that early-exit mechanism not only enhances individual performance metrics but also provides a robust and scalable solution for collaborative task offloading in large-scale UAV swarms. This approach is particularly valuable in environments where network conditions are highly dynamic, and it paves the way for future enhancements in adaptive task allocation strategies.
Figure 6: Early-exit mechanism performance as a function of the number of workers. Panel (a) shows average accuracy, (b) depicts average latency, (c) illustrates remaining GFLOPs, (d) shows the fairness index, (e) presents energy per task, and (f) shows the overall figure of merit.
VI Conclusion
In this paper, we proposed a fully distributed, diffusive metric-based solution to support split computing in large-scale UAV swarms. By introducing the concept of aggregated gigaflops, each node can dynamically compute an iterative, local measure that reflects both its own and its neighbors’ computational capacities, without requiring global synchronization or a central coordinator. This approach proved especially suitable for aerial networks with rapidly shifting topologies, heterogeneous hardware resources, and high task generation rates.
Our simulation results demonstrated that the proposed method effectively alleviates the congestion and bottlenecks that often arise in conventional centralized schemes. By shifting partial inferences (either raw inputs or intermediate layers) toward lightly loaded or computationally advantaged nodes, the swarm as a whole achieves lower latency, improved fairness in task distribution, and better energy efficiency. Moreover, we showed that our solution is robust under a variety of operational conditions, including different UAV speeds, varying mission areas, and fluctuating task arrival rates. It provides a natural mechanism for scalability: as the swarm expands and node-to-node distances change, the diffusive updates and local decisions continue to steer computational tasks toward favorable regions of the network.
The proposed approach opens several avenues for future work. One potential extension involves incorporating advanced mobility models and real flight data to capture even more realistic UAV trajectories. Additionally, while we primarily focused on split learning at the inference level, further exploration could integrate this method with federated training, enabling on-the-fly updates to model parameters in addition to distributing intermediate layers. Another intriguing direction involves dynamically adjusting the structure of the neural network layers themselves—employing model compression or pruning techniques—to further reduce communication overhead when bandwidth constraints become severe. We expect that, taken together, these refinements will push the frontier of distributed intelligence in swarms, enabling UAV networks of unprecedented scale and complexity to function reliably in even the most challenging environments.
Acknowledgment
This work is supported by The Scientific and Technological Research Council of Turkey (TUBITAK) 1515 Frontier R&D Laboratories Support Program for BTS Advanced AI Hub: BTS Autonomous Networks and Data Innovation Lab. Project 5239903 and Istanbul Technical University, Department of Scientific Research Projects (ITU-BAP, 45375), and was partially supported by the PRIN 2022 project RAIN4C: “Reliable Aerial and satellIte Networks: joint Communication, Computation, Caching for Critical scenarios” (Id: 20227N3SPN, CUP: J53D23007020001).

### v2

@@
 These results highlight the feasibility of large-scale distributed intelligence in UAV swarms and provide a blueprint for deploying robust, scalable ML services in diverse aerial networks.
-Index Terms:
+keywords:
 Split Computing , UAV Swarms , Distributed Optimization , Diffusive Metrics
-I Introduction
+1 Introduction
 The escalating prominence of Unmanned Aerial Vehicle (UAV) swarms has opened up new horizons in applications spanning surveillance, disaster response, smart agriculture, and beyond.
@@
 However, these advantages come hand-in-hand with steep technical challenges, particularly when it comes to designing and deploying machine learning (ML) services in vast swarms.
-Traditional methods that rely on a single, centralized server or controller become increasingly untenable as the number of nodes and geographic scale increase[1].
+Traditional methods that rely on a single, centralized server or controller become increasingly untenable as the number of nodes and geographic scale increase [distributed_survey_1].
 A centralized ground station, faces prohibitive communication overheads, latency bottlenecks, and reliability issues.
 Moreover, real-world UAV swarms are dynamic, meaning that nodes may join, leave, or rearrange themselves in the network.
-Such dynamic conditions only make the limitations of traditional centralized paradigms worse[2].
-Recent efforts to alleviate centralized burdens typically advocate a form of distributed operation by distributing inference or training tasks across multiple nodes[3].
+Such dynamic conditions only make the limitations of traditional centralized paradigms worse [islam2021survey].
+Recent efforts to alleviate centralized burdens typically advocate a form of distributed operation by distributing inference or training tasks across multiple nodes [distributed_survey_2].
 However, they still often rely on global awareness or explicit synchronization.
 At smaller scales or in constrained network topologies, these partial-distribution schemes might suffice.
-Yet, as UAV swarms expand to hundreds or even thousands of nodes, it becomes infeasible to gather and coordinate global network information[4].
+Yet, as UAV swarms expand to hundreds or even thousands of nodes, it becomes infeasible to gather and coordinate global network information [Tahir2019SwarmsOU].
 Additionally, heterogeneous node capabilities such as differences in on-board processing, sensing, and power resources demand an approach that can dynamically adjust to varying node states in near real-time.
@@
 These variations highlight the necessity for a self-organizing approach that pushes intelligence and decision-making to the edges.
-Split learning and computing is particularly well-suited to address these challenges in large-scale UAV swarms[3].
+Split learning and computing is particularly well-suited to address these challenges in large-scale UAV swarms [distributed_survey_2].
 Rather than transmitting raw data or entire models to a single aggregator, split learning and computing partitions the model into layers distributed among various nodes, thereby enabling partial inference at local levels.
 By transferring only the necessary, intermediate representations, nodes can reduce communication overhead while maintaining data privacy.
-However, effectively implementing split computing in UAV networks requires strategic decision-making about when and where to forward intermediate layer outputs[1, 3].
+However, effectively implementing split computing in UAV networks requires strategic decision-making about when and where to forward intermediate layer outputs [distributed_survey_1, distributed_survey_2].
 This is where centralized schemes fail to perform well for large networks.
@@
 By adjusting the exit point according to the current resource availability and task queue length, the system minimizes latency and prevents resource saturation.
-This approach aligns with the principles of early exiting [5] where intermediate outputs offer a controllable trade-off between computational effort and inference accuracy, making it particularly suitable for dynamic and resource-constrained UAV environments.
+This approach aligns with the principles of early exiting  [matsubara2023split] where intermediate outputs offer a controllable trade-off between computational effort and inference accuracy, making it particularly suitable for dynamic and resource-constrained UAV environments.
 By merging the advantages of split computing with a diffusive, graph-oriented update rule for computational metrics, our approach maintains robust inference quality across a large and rapidly evolving topology.
@@
 In this paper, our key contributions are as follows:
+  • 1.
     We identify the fundamental scalability challenges in deploying split computing across large-scale, dynamic UAV swarms, highlighting why centralized solutions are prone to failure in real-world scenarios.
+  • 2.
     We introduce a novel metric, the aggregated gigaflops, to capture how a node’s computational load should be shared in conjunction with its neighbors’ resources.
 This metric is designed to be iteratively updated at each node, enabling a fully distributed and lightweight approach.
+  • 3.
     We propose a new decision-making framework for routing partial inferences (either intermediate feature maps or raw input data) based on the aggregated gigaflops metric.
 This framework avoids reliance on global synchronization, making it highly scalable and robust to node mobility or failure.
+  • 4.
     We design an early-exit mechanism that dynamically truncates model execution based on local load conditions, reducing latency and preserving system stability under high workload scenarios.
+  • 5.
     We conduct extensive simulations to verify that our distributed solution outperforms existing benchmark strategies across several metrics including inference performance, energy efficiency, fairness, and overall throughput, particularly in large UAV swarms.
 The remainder of this paper is organized as follows.
-Section II reviews related work on distributed inference and task offloading in UAV networks.
-Section III formalizes the system model and problem definition, introducing the proposed diffusive metric and task allocation strategy.
-Section V presents the simulation setup and provides a detailed performance evaluation under various operating conditions.
-Finally, Section VI concludes the paper and outlines directions for future work.
+Section 2 reviews related work on distributed inference and task offloading in UAV networks.
+Section 3 formalizes the system model and problem definition, introducing the proposed diffusive metric and task allocation strategy.
+Section 5 presents the simulation setup and provides a detailed performance evaluation under various operating conditions.
+Finally, Section 6 concludes the paper and outlines directions for future work.
 Figure 1: Proposed distributed Split-Learning that uses diffusive aggregated Gflops.
-II Related Work
+2 Related Work
 Recent research has explored distributed DNN inference and task offloading to address the challenges posed by limited resources and dynamic network conditions.
-In [6], MoDNN partitions DNN computations among mobile devices to balance workload and reduce communication overhead; its model partitioning strategy minimizes transmission delays while exploiting parallel processing, making it highly applicable to resource-constrained environments.
-Similarly, DISNET [7] introduces a micro-split framework that partitions deep neural network models both vertically and horizontally, enabling cooperative inference among heterogeneous IoT devices while reducing latency and energy consumption through adaptive workload allocation.
-On the split computing front, [8] propose a multi tier framework that employs deep reinforcement learning to optimize model splitting and resource allocation in UE–Edge–Cloud systems, decomposing complex optimization problems into sequential subproblems to balance computational and communication loads.
-In a similar vein, CoEdge [9] demonstrates that dynamic workload partitioning—formulated as a constrained optimization problem—can yield substantial latency speedups, with a fast approximation algorithm that adapts in real time to varying device capabilities and network conditions.
-Similarly for Split Learning, [10] develop an efficient parallel split learning (EPSL) framework for resource-constrained wireless edge networks.
+In  [modnn], MoDNN partitions DNN computations among mobile devices to balance workload and reduce communication overhead; its model partitioning strategy minimizes transmission delays while exploiting parallel processing, making it highly applicable to resource-constrained environments.
+Similarly, DISNET  [disnet] introduces a micro-split framework that partitions deep neural network models both vertically and horizontally, enabling cooperative inference among heterogeneous IoT devices while reducing latency and energy consumption through adaptive workload allocation.
+On the split computing front,  [Cao2024Multitier] propose a multi tier framework that employs deep reinforcement learning to optimize model splitting and resource allocation in UE–Edge–Cloud systems, decomposing complex optimization problems into sequential subproblems to balance computational and communication loads.
+In a similar vein, CoEdge  [zeng2020] demonstrates that dynamic workload partitioning, formulated as a constrained optimization problem, can yield substantial latency speedups, with a fast approximation algorithm that adapts in real time to varying device capabilities and network conditions.
+Similarly for Split Learning,  [lin2024epsl] develop an efficient parallel split learning (EPSL) framework for resource-constrained wireless edge networks.
 Their approach jointly optimizes subchannel allocation, transmit power, and model split point selection to minimize latency and communication overhead.
-[11] propose Edge-MSL, a split learning framework for mobile edge networks that leverages a contextual multi-armed bandit formulation to dynamically select the optimal split point and assign edge servers.
+[kim2024edgemsl] propose Edge-MSL, a split learning framework for mobile edge networks that leverages a contextual multi-armed bandit formulation to dynamically select the optimal split point and assign edge servers.
 By adapting the split location and offloading decisions based on real-time network conditions, their approach significantly reduces training latency and enhances convergence.
-On the other hand, [12] combines split learning with federated learning in UAV networks to address the trade-off between communication overhead and computational constraints, resulting in an architecture that achieves higher inference accuracy than standalone federated learning while reducing transmission costs compared to conventional split learning.
-Moreover, [13] propose a parallel split learning framework over wireless networks that integrates dynamic resource management with parallel task offloading.
+On the other hand,  [hybridsplit] combines split learning with federated learning in UAV networks to address the trade-off between communication overhead and computational constraints, resulting in an architecture that achieves higher inference accuracy than standalone federated learning while reducing transmission costs compared to conventional split learning.
+Moreover,  [wu2023] propose a parallel split learning framework over wireless networks that integrates dynamic resource management with parallel task offloading.
 Their method exploits concurrent processing across heterogeneous edge devices to substantially reduce training latency while maintaining balanced workload distribution.
-[14] investigate a UAV-assisted split computing system where a UAV acts as a mobile edge server to support collaborative inference among IoT devices.
+[yeom2024uav] investigate a UAV-assisted split computing system where a UAV acts as a mobile edge server to support collaborative inference among IoT devices.
 By partitioning a DNN between on-device processing and UAV-based computation, their design reduces transmission delays and energy consumption.
 Task offloading approaches can also be useful for split learning approaches.
-For instance, [15] develops a UAV-based environmental monitoring system that employs on-board processing to minimize transmission delays and maintain low data freshness latency, thereby enhancing overall system capacity in remote areas.
-[16] proposes a cloud–MEC collaborative offloading scheme that integrates differentiated offloading decisions with service orchestration via an ODaS mechanism, significantly reducing both delay and energy consumption by balancing computational and communication loads.
-[17] addresses the challenge of offloading dependent tasks modeled as directed acyclic graphs using a branch soft actor-critic algorithm, which optimizes offloading, computation, and cooperation decisions under dynamic conditions to reduce latency and improve energy efficiency.
-More recently, [18] presents a multi-agent collaborative framework that jointly optimizes UAV trajectory planning and DAG task offloading via deep reinforcement learning, with their TD3-TT algorithm demonstrating further latency reductions and enhanced real-time performance.
-The FOCUS framework [19] leverages fog computing in UAV networks by jointly optimizing routing and computation.
+For instance,  [fei] develops a UAV-based environmental monitoring system that employs on-board processing to minimize transmission delays and maintain low data freshness latency, thereby enhancing overall system capacity in remote areas.
+[cloud_mec] proposes a cloud–MEC collaborative offloading scheme that integrates differentiated offloading decisions with service orchestration via an ODaS mechanism, significantly reducing both delay and energy consumption by balancing computational and communication loads.
+[zhou2023dag] addresses the challenge of offloading dependent tasks modeled as directed acyclic graphs using a branch soft actor-critic algorithm, which optimizes offloading, computation, and cooperation decisions under dynamic conditions to reduce latency and improve energy efficiency.
+More recently,  [zheng2024multi] presents a multi-agent collaborative framework that jointly optimizes UAV trajectory planning and DAG task offloading via deep reinforcement learning, with their TD3-TT algorithm demonstrating further latency reductions and enhanced real-time performance.
+The FOCUS framework  [focus] leverages fog computing in UAV networks by jointly optimizing routing and computation.
 Its central SDN-based architecture and analytical approach achieve significant latency improvements by effectively integrating on-board processing with data forwarding.
-Lastly, [20] investigate task offloading in a UAV-assisted MEC network by leveraging a federated learning framework.
+Lastly,  [agarwal2024federated] investigate task offloading in a UAV-assisted MEC network by leveraging a federated learning framework.
 In their work, a UAV equipped with MEC capabilities serves as a mobile server to assist resource-constrained UE in offloading computation-intensive tasks.
 Distinct from these approaches, our work targets a fully distributed split computing framework with early-exit mechanism, eliminating the need for centralized coordination in large-scale UAV swarms.
-III Problem Formulation and System Model
+3 Problem Formulation and System Model
 Figure 1 provides an overview of our proposed distributed split computing framework within a large and dynamic UAV network.
@@
 The following subsections describe each component in detail.
-III-A Task Computation Model
+3.1 Task Computation Model
 We consider machine learning tasks that are partitioned into L layers, represented as nodes in a DAG for the Machine learning task.
@@
 In scenarios where intermediate layer outputs are offloaded during processing, if the offloading decision occurs before the completion of a computation, any partially computed work is discarded.
-III-B Communications
+3.2 Communications
 We consider a UAV network in which nodes are distributed across a region and follow circular trajectories.
@@
 Lastly, $L(v_{i},v_{j})$ is the pathloss function that calculates the pathloss between nodes $v_{i}$ and $v_{j}$.
-III-C Energy Consumption Model
+3.3 Energy Consumption Model
 For energy consumption, we model that every gigaflops computed consumes $J$ joules. Therefore, the energy required to compute unit $l$ is
@@
 This expression represents the product of the transmission time $\left(\frac{S_{l}}{C_{ij}}\right)$ and the power used during transmission.
-IV Distributed Task Allocation and Early-Exit Mechanism
+4 Distributed Task Allocation and Early-Exit Mechanism
 We adopt a distributed approach for UAVs because the network spans a large geographic area with dynamic UAVs.
@@
 To identify such nodes, we aim to utilize graph centrality measures.
-However, these measures typically require a global view of the network, which conflicts with our distributed approach.
-Metrics like eigenvector centrality, which consider only a node’s neighbors, provide a feasible alternative.
-Eigenvector centrality for a node is calculated as:
-$x_{v_{i}}={\frac{1}{\lambda}}\sum_{v_{j}\in M(v_{i})}x_{v_{j}}$ (8)
-Here $M(v_{i})$ is the set of neighbors of $v_{i}$ and $\lambda$ is the largest eigenvalue of the network graph’s adjacency matrix.
-Eigenvector centrality is iteratively computed using power iteration: initial centralities are guessed and updated until convergence.
-Normalization is required at each step to prevent divergence, though this typically requires a global view.
-This iterative approach could work for distributed networks if divergence can be controlled.
-To address these challenges, we propose using aggregated computation capability instead of eigenvector centrality.
-This metric measures how quickly a node can complete a job when it distributes the load among itself and its neighbors.
+Metrics like eigenvector centrality, which propagate importance through neighbors, appear suitable for capturing this effect.
+However, computing eigenvector centrality requires global normalization at every iteration, making it impractical for a fully distributed and rapidly changing UAV network.
 Figure 2: Proposed early-exit mechanism
-We iteratively compute aggregated computation capability.
-The idea is if a load given to a node, if it were to distribute this load to its neighbors evenly, when this load would be completed.
-The aggregated computation capability $\phi_{i,t+1}$ of node $v_{i}$ at time $t+1$ is:
+To address this issue, we propose using aggregated computation capability instead of eigenvector centrality.
+This metric preserves the neighbor-based recursive structure while removing the need for global normalization.
+Eigenvector centrality updates each node using the contributions of its neighbors, and our formulation mirrors this idea: each node’s effective computation capability is updated based on its own resources and the aggregated capability of its neighbors.
+We compute aggregated computation capability iteratively.
+The intuition is that when a node splits an incoming load evenly among itself and its neighbors, the total completion time is dictated by the slowest participant.
+This leads to the following neighbor-recursive update for the aggregated computation capability $\phi_{i,t+1}$ of node $v_{i}$ at time $t+1$:
 $\frac{1}{\phi_{i,t+1}}=\frac{1}{|M_{i}(t)|+1}\big(\frac{1}{F_{i}}+\max_{k\in M_{i}(t)}\{d^{tx}_{i,k}+\frac{1}{{\phi_{k,t}}}\}\big)$ (8)
-where $M_{i}(t)$ is the neighbor set of node $i$, $F_{i}$ is computation capability of node $i$, $d^{tx}_{i,k}$ is the transmission delay between nodes $i$ and $k$, and $\phi_{i,t}$ is the aggregated computation capability of node $i$ and time $t$.
-The aggregated computation capability assumes that the node divides the work evenly ($|M(i)|+1$), computes its part, then it broadcasts the load to its neighbors.
-Since the neighbors can work concurrently, the task will be completed lastly by the slowest node with aggregated computation capability.
-Since this update requires only the neighbors of the nodes, it dynamically adapts the network conditions without need for global view.
-IV-A Task Transfer Decision
+where $M_{i}(t)$ is the neighbor set of node $i$, $F_{i}$ is computation capability of node $i$, $d^{tx}_{i,k}(t)$ is the transmission delay between nodes $i$ and $k$, and $\phi_{i,t}$ is the aggregated computation capability of node $i$ and time $t$.
+The aggregated computation capability assumes that a given amount of work is divided evenly among the node and its neighbors ($|M_{i}(t)|+1$).
+The node processes its own share and forwards the remaining portions to its neighbors.
+Since all participants work concurrently, the overall completion time is determined by the slowest node, whose aggregated capability is captured by the maximum term.
+Because this update requires only the neighbors of the nodes, it naturally adapts to dynamic network conditions without requiring any global view.
+4.1 Task Transfer Decision
 Our framework requires each UAV to make autonomous decisions about whether to process a task locally or transfer it to a neighboring node.
@@
 This mechanism ensures that tasks are only transferred when the utilization difference exceeds a meaningful threshold, preventing oscillatory behavior where tasks might be repeatedly transferred between nodes with similar utilization levels.
-IV-B Early-Exit Mechanism
+4.2 Early-Exit Mechanism
 To maintain consistent performance under varying workloads, nodes dynamically adjust the depth of the model inference based on real-time buffer conditions.
@@
 This adaptive mechanism ensures that nodes under high load degrade computation gracefully, trading off accuracy to maintain responsiveness and system stability.
-It is important to note that the neural network architecture, the selected split points, and the number of layers used in this work serve as a specific example corresponding to the evaluation presented in Section V.
+It is important to note that the neural network architecture, the selected split points, and the number of layers used in this work serve as a specific example corresponding to the evaluation presented in Section 5.
 In practice, the neural network designer must analyze the target model to identify appropriate split points and determine the number of additional layers to execute after each truncation.
 These decisions depend on the network’s structure and the desired balance between accuracy and computational efficiency under dynamic conditions.
-V Performance Evaluation
-Our simulation environment is designed to emulate the dynamic conditions of a large-scale UAV swarm executing split computing tasks.
-In our environment, UAVs are deployed over a 20×20 km area and follow circular trajectories with a movement radius of 1000 m, traveling at speeds of up to 75 m/s.
-Task arrivals are modeled as a Markov process with an average inter-arrival time of 60 ms, which accurately captures the random and bursty nature of workload generation.
-In addition, the simulator captures detailed communication and computation aspects—including realistic channel characteristics, transmission delays, and the heterogeneous computational capabilities of each UAV ensuring that both the physical movement and the fluctuating workload distribution among nodes are properly considered.
-Table I shows the rest of the simulation parameters.
-Within this dynamic simulation framework, we evaluate five distinct task offloading strategies.
-In the first scenario, tasks are offloaded in a purely random manner to a neighboring UAV, providing a basic baseline for performance comparison.
-The second strategy follows a similar random selection process, but enforces an acyclic constraint to prevent redundant offloading cycles.
-The third approach adopts a greedy policy, where each UAV offloads its task to the least busy neighbor based on instantaneous load measurements, thus aiming to minimize processing delays.
-The fourth strategy, referred to as LocalOnly, eliminates offloading entirely so that tasks are processed solely on the originating UAV, which highlights the trade-off between local processing overhead and the benefits of collaborative offloading.
-Finally, our proposed Distributed method leverages a diffusive metric—in particular, the aggregated gigaflops measure—to dynamically route tasks based on both the local UAV’s capacity and the capabilities of its neighbors.
-This comprehensive set of cases allows us to rigorously assess how different offloading mechanisms perform under the challenging conditions of a mobile, Markov-driven UAV swarm.
-In our experiments, we consider the neural network model described in Section III pre-trained offline and deployed across the UAV swarm.
-The full model achieves an accuracy of 0.95, while the early-exit configurations at the 30th and 15th layers provide accuracies of 0.9 and 0.6, respectively.
-These values quantify the trade-off between computational savings and inference quality when applying early exits under high-load conditions.
-Lastly, the threshold $\gamma$ for distributed allocation is set to 0.02 based on our empirical data gathered from the empirical simulations.
-We selected this value to effectively balance the trade-off between frequent task transfers and achieving equitable workload distribution.
-TABLE I: Simulation Parameters
+Table 1: Simulation Parameters
 Parameter Value
@@
 Greedy Allocation Probability 0.05
-V-A Basic Simulations
+5 Performance Evaluation
+Our simulation environment is designed to emulate the dynamic conditions of a large-scale UAV swarm executing split computing tasks.
+In our environment, UAVs are deployed over a 20×20 km area and follow circular trajectories with a movement radius of 1000 m, traveling at speeds of up to 75 m/s.
+Task arrivals are modeled as a Markov process with an average inter-arrival time of 60 ms, which accurately captures the random and bursty nature of workload generation.
+In addition, the simulator captures detailed communication and computation aspects,including realistic channel characteristics, transmission delays, and the heterogeneous computational capabilities of each UAV ensuring that both the physical movement and the fluctuating workload distribution among nodes are properly considered.
+Table 1 shows the rest of the simulation parameters.
+Within this dynamic simulation framework, we evaluate five distinct task offloading strategies.
+In the first scenario, tasks are offloaded in a purely random manner to a neighboring UAV, providing a basic baseline for performance comparison.
+The second strategy follows a similar random selection process, but enforces an acyclic constraint to prevent redundant offloading cycles.
+The third approach adopts a greedy policy, where each UAV offloads its task to the least busy neighbor based on instantaneous load measurements, thus aiming to minimize processing delays.
+The fourth strategy, referred to as LocalOnly, eliminates offloading entirely so that tasks are processed solely on the originating UAV, which highlights the trade-off between local processing overhead and the benefits of collaborative offloading.
+Finally, our proposed Distributed method leverages a diffusive metric, to dynamically route tasks based on both the local UAV’s capacity and the capabilities of its neighbors.
+This comprehensive set of cases allows us to rigorously assess how different offloading mechanisms perform under the challenging conditions of a mobile, Markov-driven UAV swarm.
+In our experiments, we consider the neural network model described in Section 3 pre-trained offline and deployed across the UAV swarm.
+The full model achieves an accuracy of 0.95, while the early-exit configurations at the 30th and 15th layers provide accuracies of 0.9 and 0.6, respectively.
+These values quantify the trade-off between computational savings and inference quality when applying early exits under high-load conditions.
+Lastly, the threshold $\gamma$ for distributed allocation is set to 0.02 based on our empirical data gathered from the empirical simulations.
+We selected this value to effectively balance the trade-off between frequent task transfers and achieving equitable workload distribution.
+Figure 3: Simulation results as a function of the number of workers.
+(a) shows average latency, (b) depicts remaining GFLOPs, (c) illustrates average transfer time, (d) presents Jain’s fairness index, (e) reports energy per task, and (f) shows the overall figure of merit.
+5.1 Basic Simulations
 The initial set of simulations examines a baseline scenario that does not employ early-exit mechanism.
@@
 These findings underscore the capability of our system to efficiently manage computational loads in dynamic UAV environments without centralized control, while ensuring both high resource utilization and robustness against congestion.
-Figure 3: Simulation results as a function of the number of workers.
-(a) shows average latency, (b) depicts remaining GFLOPs, (c) illustrates average transfer time, (d) presents Jain’s fairness index, (e) reports energy per task, and (f) shows the overall figure of merit.
-V-B Effect of Input Rate
+Figure 4: Performance metrics under varying task arrival rates for 30 workers. Figure (a) shows average latency, (b) displays remaining GFLOPs, and (c) presents the overall figure of merit.
+Figure 5: Performance metrics under varying mission area sizes. Figure (a) shows average latency, (b) depicts remaining GFLOPs, and (c) presents the overall figure of merit.
+5.2 Effect of Input Rate
 Next, we measure the effect of input rate when there is fixed number of workers.
@@
 This makes sense because when there are not many tasks, the overhead of collaboration does provide significant benefits.
-Figure 4: Performance metrics under varying task arrival rates for 30 workers. Panel (a) shows average latency, (b) displays remaining GFLOPs, and (c) presents the overall figure of merit.
-V-C Effect of Network Distribution
+5.3 Effect of Network Distribution
 The performance of collaborative task offloading is further influenced by the spatial distribution of UAVs, as captured by variations in the mission area size.
@@
 These observations suggest that for large-scale deployments, additional strategies to enhance connectivity or adaptive routing protocols may be necessary to sustain optimal performance.
-Figure 5: Performance metrics under varying mission area sizes. Panel (a) shows average latency, (b) depicts remaining GFLOPs, and (c) presents the overall figure of merit.
-V-D Early-Exit Mechanism
+Figure 6: Early-exit mechanism performance as a function of the number of workers.
+Figure (a) shows average accuracy, (b) depicts average latency, (c) illustrates remaining GFLOPs, (d) shows the fairness index, (e) presents energy per task, and (f) shows the overall figure of merit.
+5.4 Early-Exit Mechanism
 In the final set of experiments, we evaluate the efficacy of our early-exit mechanism, which adapts the network configuration in real-time based on load and buffer conditions executing an early-exit mechanism.
@@
 This approach is particularly valuable in environments where network conditions are highly dynamic, and it paves the way for future enhancements in adaptive task allocation strategies.
-Figure 6: Early-exit mechanism performance as a function of the number of workers.
-Panel (a) shows average accuracy, (b) depicts average latency, (c) illustrates remaining GFLOPs, (d) shows the fairness index, (e) presents energy per task, and (f) shows the overall figure of merit.
-VI Conclusion
+6 Conclusion
 In this paper, we proposed a fully distributed, diffusive metric-based solution to support split computing in large-scale UAV swarms.
@@
 Additionally, while we primarily focused on split learning at the inference level, further exploration could integrate this method with federated training, enabling on-the-fly updates to model parameters in addition to distributing intermediate layers.
-Another intriguing direction involves dynamically adjusting the structure of the neural network layers themselves—employing model compression or pruning techniques—to further reduce communication overhead when bandwidth constraints become severe.
+Another intriguing direction involves dynamically adjusting the structure of the neural network layers themselves, employing model compression or pruning techniques, to further reduce communication overhead when bandwidth constraints become severe.
 We expect that, taken together, these refinements will push the frontier of distributed intelligence in swarms, enabling UAV networks of unprecedented scale and complexity to function reliably in even the most challenging environments.
 Acknowledgment
-This work is supported by The Scientific and Technological Research Council of Turkey (TUBITAK) 1515 Frontier R&D Laboratories Support Program for BTS Advanced AI Hub: BTS Autonomous Networks and Data Innovation Lab.
-Project 5239903 and Istanbul Technical University, Department of Scientific Research Projects (ITU-BAP, 45375), and was partially supported by the PRIN 2022 project RAIN4C: “Reliable Aerial and satellIte Networks: joint Communication, Computation, Caching for Critical scenarios” (Id: 20227N3SPN, CUP: J53D23007020001).
+This work is supported by Istanbul Technical University, Department of Scientific Research Projects (ITU-BAP, 45375), and was partially supported by the PRIN 2022 project RAIN4C: “Reliable Aerial and satellIte Networks: joint Communication, Computation, Caching for Critical scenarios” (Id: 20227N3SPN, CUP: J53D23007020001).